# Boosting Latent Diffusion with Flow Matching

### v1

@@ 3.2.3 Noise Augmentation
Noise Augmentation is a technique for boosting generative models’ performance introduced for cascaded Diffusion models [19]. The authors found that applying random Gaussian noise or Gaussian blur to the conditioning signal in super-resolution Diffusion models results in higher-quality results during inference. Drawing inspiration from this, we also implement Gaussian noise augmentation on $x_{0}$. Following the notation from variance-preserving DMs, we noise $x_{0}$ according to the cosine schedule first proposed in [43]. In line with [19], we empirically discover in our experiments that incorporating a specific amount of Gaussian noise enhances performance. We hypothesize that including a small amount of Gaussian noise smoothes the base probability density $p_{0}$ so that it remains well-defined over the higher-dimensional space. Note that this noise augmentation is only applied to $x_{0}$ but not to the conditioning information $z$, since the model relies on the precise conditioning information to construct the straight path.
Figure 4: Uncurated samples from the Coupling Flow Matching model on top of Stable Diffusion V1.5 [48] using a classifier-free guidance scale of $7.5$. Samples are generated in latent space $64^{2}$ and up-sampled with CFM from $64^{2}$ to $128^{2}$. The resulting images have a resolution of $1024\times 1024$.
@@ 3.2.4 Latent Flow Matching
In order to reduce the computational demands associated with training Flow Matching models for high-resolution image synthesis, we take inspiration from [48, 12] and utilize an autoencoder model that provides a compressed latent space that aligns perceptually with the image pixel space similar to LDMs. By training in the latent space, we get a three-fold advantage: i) The computational cost associated with the training of Flow Matching models is reduced substantially, thereby enhancing the overall efficiency of the training process. ii) Leveraging the latent space unlocks the potential to synthesize images with significantly increased resolution efficiently. iii) The lower resolution in the latent space also yields quicker inference speed in the synthesis task.
@@ 3.3 High-Resolution Image Synthesis
Overall, our approach integrates all the components discussed above into a cohesive synthesis pipeline, as depicted in detail in Fig. 3. We start from a DM for content synthesis and move the generation to a latent space with a pretrained VAE encoder, which optimizes memory usage and enhances inference speed. To further alleviate the computational load of DM and achieve additional acceleration, we adopt a relatively compact DM that produces compressed information. Subsequently, the FM model projects the compressed information to a high-resolution latent image with a straight conditional probability path. Finally, we decompress the latent space using a pretrained VAE decoder. Note that the VAE decoder performs well across various resolutions, as we demonstrate in Fig. 16 and Tab. 6 in the supplementary materials.
The integration of FM with DMs in the latent space presents a promising approach to address the trade-off between flexibility and efficiency in modeling the dynamic image synthesis process. The inherent stochasticity within a DM’s sampling process allows for a more nuanced representation of complex phenomena, while the FM model exhibits greater computational efficiency, which is useful when handling high-resolution images, but lower flexibility and image fidelity as of yet when it comes to image synthesis [32]. By combining them in the pipeline, we benefit from the flexibility of the DM while capitalizing on the efficiency of FM as well as VAE.
@@ 4 Experiments
Combining a diffusion synthesis model and a flow matching super-resolution model yields great synthesis capability in the high-resolution domain, as we show qualitatively and quantitatively in this section. We first demonstrate the effectiveness of the combination, and then provide an extensive ablation and analysis of our CFM model. Subsequently, we assess our model in comparison to the current state-of-the-art diffusion approach.
4.1 Datasets
The general dataset that we use for training and ablations is FacesHQ, a compilation of CelebA-HQ [24] and FFHQ [25], as employed in previous work such as [13]. This dataset serves as a foundation for evaluating various aspects of our model. We also include ablations on LHQ [55], which contains 90k high-resolution landscape images.
For the general flow matching (FM) model, particularly targeted towards combination with latent diffusion models, we utilize the Unsplash dataset [3], which provides diverse and high-quality images for training our model. We present inference results of our general FM model on a high-resolution subset of LAION 5B [53].
4.2 Boosting LDM with FM
Aggregating the power of LDM and FM results in remarkable performance in image synthesis. Combining the flexibility of LDM with the intrinsic characteristics of FM achieves an optimal trade-off between computational efficiency and visual fidelity.
Since LDMs are typically trained and evaluated on fixed sizes, we follow [22] and apply attention scaling on SD to synthesize images of varying resolutions. We also finetune the small LDM model which outputs images at resolutions of $256^{2}$ px to further ensure image fidelity. More details can be found in Appendix Sec. 6.2. We plot the inference speed for $1024^{2}$ px image in Fig. 1, where we visualize the time taken by LDM and FM distinctively. Note that the LDM’s inference time scales quadratically with increasing resolution, and the inference is almost impractical for a real-time inference for a latent space of $128^{2}$. Conversely, the CFM model shows remarkably fast inference speed, which emphasises the efficiency improvements achieved by its integration for high-resolution image synthesis.
We demonstrate the models’ synthesized image quality in Fig. 5. Both combinations of LDM and FM surpass the performance of the attention-scaled LDM-128 baseline in terms of inference time and image quality. The intermediate resolution of $32^{2}$ yields faster inference, while the intermediate resolution of $64^{2}$ maintains better image fidelity. Note that we achieve near-optimal results with as few as $10$ ODE steps for the CFM model, which adds minimal computational overhead to the underlying DM. Note that we selectively use a subset of $1$k samples from LAION to highlight variations in NFE for a fair comparison across all models. This choice is for rapid evaluation purposes, and the FID is expected to decrease with a larger sample size. We present a selection of image samples of LDM-64 and CFM $64^{2}\rightarrow 128^{2}$ in Fig. 4. Please refer to Fig. 11 and Fig. 14 in the Appendix for more samples.
We also investigate the results for consistency models [59], which significantly accelerate the inference process by distilling information from a diffusion model. The LCM-LoRA model [39] finetuned on SD V1.5 [48] produces $512\times 512$ resolution images. We can further improve this model to 1k resolution images with our CFM model. We present our synthesised results in Fig. 2. The inference time for a batch of four samples is $1.388$ seconds. The LCM-LoRA SDXL model [44] fails to produce images with similar fidelity at the same resolution in the same time range.
Figure 5: Inference time per model part (LDM or CFM) and FID for synthesizing images with resolution $1024\times 1024$ for different number of function evaluations (NFE). The upper plot refers to the average inference time relative to LDM-128, while the lower plot indicates $\log$ FID for $1$k samples. dopri5 corresponds to the Dormand-Prince adaptive step-size solver. For LDM-128 we do not use any CFM model.
4.3 FM Model Ablation
Upsampling Methods As our flow matching model requires that $x_{0}$ and $x_{1}$ share the same dimensionality, we need to upsample the latent code. In this context, we conduct an ablation study comparing two upsampling methods: nearest neighbor and bi-linear upsampling. The results are presented in Tab. 1. While the choice between these methods does not appear to be decisive based on SSIM and PSNR metrics, we observe generally lower FID [16] values with nearest neighbor upsampling.
CelebA-HQ FFHQ FacesHQ
Model SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
CFM-nearest 0.75 26.47 4.00 0.70 25.47 3.65 0.72 25.80 2.75
CFM-bilinear 0.75 26.43 4.58 0.70 25.50 3.94 0.70 25.49 3.93
FM-nearest 0.75 26.47 7.91 0.70 25.47 8.16 0.71 25.80 6.84
FM-bilinear 0.75 26.47 8.10 0.69 25.47 7.98 0.71 25.80 6.78
Table 1: Ablation of up-sampling methods for Flow Matching in the latent space. While the choice between these methods does not appear to be decisive based on SSIM and PSNR metrics, we observe generally lower FID values with nearest neighbor upsampling which also matches the visual inspection of generated results shown in the supplementary.
LR
Reg
FM
CFM [Uncaptioned image]
Figure 6: Results of different latent space up-sampling strategies from $32\times 32$ to $128\times 128$. LR corresponds to bilinear upsampling of the low-resolution image, Reg refers to the regression baseline. FM and CFM correspond to Flow Matching and Coupling Flow Matching with a noising step of 400, respectively.
GT REG LR FM
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 7: Qualitative comparison of the Flow Matching (FM) model to the upsampled low resolution image (LR) the regression model (REG) and the ground truth (GT) using zoomed-in patches. The FM patches are significantly sharper and display more fine-grained structure and details.
CelebA-HQ FFHQ FacesHQ
Model SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
Reg 0.81 26.99 56.15 0.77 26.29 48.22 0.78 26.52 48.32
FM 0.75 26.47 07.91 0.70 25.47 08.16 0.71 25.80 06.84
CFM 0.75 26.47 04.00 0.70 25.47 03.65 0.72 25.80 02.75
Table 2: Metric results for regression (Reg), Flow Matching (FM), and Coupling Flow Matching (CFM) corresponding to Fig. 6.
Flow Matching and Coupling Flow Matching In this ablation study, we systematically explore two variations of Flow Matching models – one incorporating the Gaussian assumption (FM) and the other incorporating data-dependent coupling (CFM). We evaluated these two variants quantitatively (Tab. 1) and qualitatively (Fig. 7), where we observed that the CFM models with data-dependent coupling readily outperform the ones without. We provide more information about the noise augmentation process in Tab. 3. Notably, in the specific upsampling scenario from $256^{2}$ to $1024^{2}$, we observe an optimal configuration with a noising timestep of 200. The introduction of Gaussian noise proves beneficial as it imparts a smoothing effect on the input probability path, resulting in improved performance. However, excessive Gaussian noise can lead to the loss of valuable information, subsequently deteriorating the data-dependent coupling and reverting the model’s behavior to the Gaussian assumption. This finding underscores the delicate balance required in incorporating noise for optimal model performance.
Intermediate Results along the ODE Trajectory Fig. 8 shows intermediate results along the ODE trajectory. It can be seen that the CFM model gradually transforms the image representation from one with grid artefacts due to nearest neighbour upsampling to its high-resolution image counterpart.
CelebA-HQ FFHQ FacesHQ
Model SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ FID $\downarrow$
CFM-100 0.71 25.30 8.27 0.73 26.14 5.84 0.75 26.41 5.47
CFM-200 0.75 26.39 3.90 0.70 25.48 3.14 0.72 25.78 2.38
CFM-400 0.75 26.47 4.00 0.70 25.47 3.65 0.72 25.80 2.75
CFM-600 0.75 26.51 5.18 0.70 25.52 5.14 0.72 25.85 4.04
CFM-999 0.74 26.36 7.72 0.69 25.35 7.53 0.71 25.69 6.38
Table 3: Quantative evaluation on different noising steps for coupling flow matching (CFM) for 4×4$\times$4 × upsampling from $256^{2}$ to $1024^{2}$ px. We evaluate the FID on the full dataset (5k, 10k, and 15k samples respectively).
@@ 4.3 FM Model Ablation
Figure 9: Comparison of diffusion-based SR3 [51] and our Coupling Flow Matching (CFM) over the training for 4×4$\times$4 × up-sampling of the latent codes from $32^{2}\rightarrow 128^{2}$. Decoded output resolution is $1024^{2}$. Architecture and hyperparameters are kept fixed for both. FID evaluated on $1$k samples from the validation set. We use DDIM sampling with $50$ steps for the diffusion-based model and the Euler ODE solver with $50$ steps for CFM.
Figure 10: Comparison of diffusion-based SR3 [51] and our Coupling Flow Matching (CFM) for different number of function evaluations (NFE) for 4×4$\times$4 × up-sampling of the latent codes from $32^{2}\rightarrow 128^{2}$. Decoded output resolution is $1024^{2}$. Architecture and hyperparameters are kept fixed for both. FID evaluated on $5$k samples from the validation set. We use DDIM [57] for diffusion-based sampling and the Euler ODE solver for the CFM model.
4.4 Comparison to Baselines
We compare our Coupling Flow Matching method to the diffusion-based up-sampling method SR3 [51], as well as a simple regression baseline, trained with a simplep $L2$ loss, on the FacesHQ dataset. For a fair comparison, we keep the architecture and hyperparameters fixed for all three models. We obtain the latent codes by using the pre-trained KL-regularized autoencoder from [48] and perform nearest neighbor up-sampling to fit the dimensionality of our starting condition $x_{0}$.
Upsampling Results
Fig. 6 shows results of different latent space up-sampling strategies from $32^{2}$ to $128^{2}$, increasing the final output resolution from $256^{2}$ to $1024^{2}$. We can observe that the simple regression baseline yields blurry results, which is also reflected in the metrics (see Tab. 2). In contrast, both flow matching approaches – Coupling Flow Matching (CFM) and Flow Matching starting from a noise distribution (FM) – yield better results. Also, metric-wise, both outperform the simple regression baseline by a large margin. When directly comparing the low-resolution image at size $256^{2}$ with the decoded output of our Flow Matching model (Fig. 6), we can observe that FM adds fine-grained details to the image not visible in the lower-resolution counterpart.
Training Efficiency
@@ 4.4 Comparison to Baselines
Based on optimal transport theory, the training of a constant velocity field presents a more straightforward training objective when contrasted with the intricate high-curvature probability paths found in diffusion models [32, 12]. This distinction often translates to slower training convergence and potentially sub-optimal trajectories for DMs, which could detrimentally impact both, training duration and overall model performance. Fig. 9 shows the FID over the course of training for the diffusion-based SR3 and our Flow Matching model. We can clearly observe that the Flow Matching model achieves a lower FID more rapidly and produces superior results, compared to the SR3 model. Fig. 10 further demonstrates that after $100$k iterations and for different number of function evaluations (NFE), we consistently attain a lower FID compared to the diffusion-based SR3 model. Taken together, these findings underscore the training efficiency of Flow Matching over Diffusion models and its superior performance for the up-sampling task.

### v2

@@ 3.2.3 Noise Augmentation
Noise Augmentation is a technique for boosting generative models’ performance introduced for cascaded Diffusion models [20]. The authors found that applying random Gaussian noise or Gaussian blur to the conditioning signal in super-resolution Diffusion models results in higher-quality results during inference. Drawing inspiration from this, we also implement Gaussian noise augmentation on $x_{0}$. Following variance-preserving DMs, we noise $x_{0}$ according to the cosine schedule first proposed in [47]. In line with [20], we empirically discover that incorporating a specific amount of Gaussian noise enhances performance. We hypothesize that including a small amount of Gaussian noise smoothes the base probability density $p_{0}$ so that it remains well-defined over the higher-dimensional space. Note that this noise augmentation is only applied to $x_{0}$ but not to the conditioning information $z$, since the model relies on the precise conditioning information to construct the straight path.
@@ 3.2.4 Latent Flow Matching
In order to reduce the computational demands associated with training FM models for high-resolution image synthesis, we take inspiration from [53, 13] and utilize an autoencoder model that provides a compressed latent space that aligns perceptually with the image pixel space similar to LDMs. By training in the latent space, we get a two-fold advantage: i) The computational cost associated with the training of flow-matching models is reduced substantially, thereby enhancing the overall training efficiency. ii) Leveraging the latent space unlocks the potential to synthesize images with significantly increased resolution efficiently and with a faster inference speed.
@@ 3.3 High-Resolution Image Synthesis
Overall, our approach integrates all the components discussed above into a cohesive synthesis pipeline, as depicted in detail in Fig. 2. We start from a DM for content synthesis and move the generation to a latent space with a pretrained VAE encoder, which optimizes memory usage and enhances inference speed. To further alleviate the computational load of the DM and achieve additional acceleration, we adopt a relatively compact DM that produces compressed information. Subsequently, the FM model projects the compressed information to a high-resolution latent image with a straight conditional probability path. Finally, we decompress the latent space using a pre-trained VAE decoder. Note that the VAE decoder performs well across various resolutions, we show further proof in the appendix.
The integration of FM with DMs in the latent space presents a promising approach to address the trade-off between flexibility and efficiency in modeling the dynamic image synthesis process. The inherent stochasticity within a DM’s sampling process allows for a more nuanced representation of complex phenomena, while the FM model exhibits greater computational efficiency, which is useful when handling high-resolution images, but lower flexibility and image fidelity as of yet when it comes to image synthesis [36]. By combining them in the pipeline, we benefit from the flexibility of the DM while capitalizing on the efficiency of FM as well as a VAE.
@@ 4 Experiments
Figure 4: Uncurated samples from the Coupling Flow Matching model on top of SD 1.5 [53] using a classifier-free guidance scale of $7.5$. Samples are generated in latent space $64^{2}$ and up-sampled with CFM from $64^{2}$ to $128^{2}$. The resulting images have a resolution of $1024\times 1024$ pixels. Best viewed via zoomed in.
4.1 Metrics and datasets
For quantitative evaluation, we use the standard Fréchet Inception Distance (FID)[17], SSIM[70], and PSNR to measure the realism of the output distribution and the quality of the image. The general dataset we use for initial experiments and ablations is FacesHQ, a compilation of CelebA-HQ [25] and FFHQ [26], as used in previous work [14, 56] for high-resolution synthesis tasks. However, as highlighted in [11], FID struggles to capture detail and measure fidelity at higher resolutions. To remedy this, we also report p-FID [17] for a more comprehensive evaluation, especially when images contain objects at different scales, such as LHQ [60], which contains 90k high-resolution landscape images and offers a more diverse scale of scenes/objects presented in the image compared to FacesHQ. These two datasets serve as the basis for the evaluation.
For the general T2I image synthesis task, we train on the Unsplash dataset [2], which provides diverse and high-quality images for training our model. Later, we evaluate on a high-resolution subset of LAION-5B [58] to check how well the model generalizes to unseen data.
4.2 Boosting LDM with CFM
Zero-shot LAION-5k 1024$\times$1024
Model CLIP $\uparrow$ FID $\downarrow$ p-FID $\downarrow$ time (s/im) $\downarrow$
SD1.5 + CFM $\bf{256^{2}\rightarrow 1024^{2}}$ 23.75 25.47 23.31 0.62
SD1.5 + CFM $\bf{512^{2}\rightarrow 1024^{2}}$ 26.14 21.67 15.96 3.16
LCM-LoRA SDXL [43] 24.51 28.98 24.00 1.83
Table 1: Quantitative comparison for $1024^{2}$ image synthesis using SD v1.5 [53] plus our Coupling Flow Matching (CFM) method against a state-of-the-art diffusion speed-up method. The numbers after CFM symbolize the starting and ending resolutions in pixel space. FID and Patch-FID are computed for 5k samples. We use the fixed step-size Euler ODE solver with $40$ number of function evaluations for CFM. For LCM-LoRA SDXL [43] we use $4$ sampling steps.
Combining LDM with CFM achieves an optimal trade-off between computational efficiency and visual fidelity. We visualize the time taken by LDM and FM, respectively, to synthesize 1k resolution images in Fig. 5, where LDM’s inference time scales quadratically with increasing resolution, and inference is nearly impractical for real-time inference for a latent space of $128^{2}$. To ensure a fair comparison within the limited time frame, we compare our combination to the LCM-LoRA SDXL model [64, 43], which is known for its significantly faster inference than the original SDXL model. Tab. 1 shows that our approach with a standard SD baseline model yields superior performance in terms of FID and inference speed. Note that we apply attention scaling [23] on SD to synthesize images for varying resolutions and finetune the models, with more details in the Appendix. We present a selection of image samples from the baseline SD1.5 model and CFM $64^{2}\rightarrow 128^{2}$ in Fig. 4. We can equally upscale the LCM-LoRA SD1.5 model from 512 to 1k resolution images with our CFM model. We present our synthesized results in Fig. 1. The inference time for a batch of four samples is $1.388$ seconds on an NVIDIA A100 GPU. The LCM-LoRA SDXL model fails to produce images with similar fidelity at the same resolution within the same time.
We further demonstrate the effectiveness of our approach by comparing it to state-of-the-art models [76, 48, 49] in image synthesis on COCO $1024\times 1024$, including CogView3 [76]. We reduce the computational cost of the diffusion component by using a lower resolution and fewer steps, while offloading the remaining steps to our CFM module. This approach significantly reduces the inference time and maintains a good trade-off between speed and accuracy, as shown by the FID in Tab. 2. In summary, we achieve a competitive FID at a faster inference speed than the counterpart diffusion models.
Zero-shot COCO-5k 1024$\times$1024
@@ 4.2 Boosting LDM with CFM
    SDXL [49] 50 19.67s 26.29
    StableCascade [48] 20+10 10.83s 36.59
    CogView3 [76] 50+10 10.33s 31.63
$\bf{256^{2}\rightarrow 1024^{2}}$
    SD1.5 + CFM 25+20 4.07s 33.48
    SD1.5 + CFM 40+40 5.88s 30.64
$\bf{512^{2}\rightarrow 1024^{2}}$
    SD1.5 + CFM 25+20 8.79s 28.81
Table 2: Metric results on 5k samples from the COCO dataset [35]. All samples are generated on $1024\times 1024$. The time cost is measured with a batch size of 4. Table adapted from [76].
Figure 5: Comparison of 1k image synthesis performance using different architectures. We utilize SD v1.5 as our base model for LDM and adapt its resolution based on [23]. LDM’s inference time grows quadratically with higher resolutions, making real-time inference nearly impractical at a $128^{2}$ resolution latent space. In contrast, the integration of Coupling Flow Matching (CFM) with $50$ function evaluations exhibits consistently faster inference, highlighting its efficiency in high-resolution image synthesis.
Figure 6: Samples synthesized in $2048^{2}$ px. The base diffusion model is SD1.5 synthesizing images in $512^{2}$ px. By plugging in our CFM module, we can quickly boost the resolution to 2k and generate high-fidelity images. Best viewed when zoomed in.
GT NFE=1 NFE=2 NFE=4 NFE=10 NFE=50
Figure 7: Sample quality for different number of function evaluations (NFE). From left to right, 1st column represents the ground truth, high-resolution image. From the 2nd column on, we show the results for NFE $={1,2,4,10,50}$ with the Euler ODE solver.
4.3 Baseline Comparison
We compare our CFM model to three baseline methods on the FacesHQ and LHQ datasets. For a fair comparison, we fix the UNet architecture and hyperparameters so that the models only differ in their respective training objectives.
Low-Res L2 FM CFM
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Results for different baseline methods, increasing resolution from $256^{2}$ px to $1024^{2}$ px. Low-Res corresponds to bi-linear upsampling of the low-resolution image, L2 refers to the L2 regression baseline. FM and CFM correspond to Flow Matching and Coupling Flow Matching, respectively. Best viewed when zoomed in.
Regression Baseline. Similar to [56], we compare simple one-step regression models with $L1$ and $L2$ loss, respectively. The input is the low-resolution latent code and the target is the corresponding high-resolution latent code of the pre-trained KL autoencoder. In contrast, our method is trained with $L2$ loss on intermediate vector fields. Tab. 3 shows that CFM yields superior metric results. This is also reflected qualitatively, as visualized in Fig. 8, where the images from the regression baseline are visually blurry due to the mode-averaging behavior of the MSE regression. CFM excels at adding fine-grained, high-resolution detail to the image. We conclude that simple regression models trained with $L1$ or $L2$ loss are not sufficient to increase resolution in latent space.
FacesHQ LHQ
[…]
$L1$ 0.86 31.78 4.52 6.51 0.72 26.99 4.88 6.54
$L2$ 0.85 31.48 5.73 9.07 0.72 26.87 6.02 8.59
DM 0.73 23.68 2.72 4.71 0.61 19.94 4.29 4.55
FM 0.82 30.46 1.37 2.10 0.68 25.50 2.31 2.61
CFM (ours) 0.82 30.40 1.36 1.62 0.69 25.69 2.27 2.38
Table 3: Metric results for $L1$ and $L2$ regression, diffusion-based (DM) similar to [56], Flow Matching (FM) [36], and our Coupling Flow Matching (CFM) on 5k samples from FacesHQ and LHQ high-resolution datasets, respectively.
@@ 4.3 Baseline Comparison
Diffusion Models. Based on optimal transport theory, the training of a constant velocity field presents a more straightforward training objective when contrasted with the intricate high-curvature probability paths found in DMs [36, 13]. This distinction often translates to slower training convergence and potentially sub-optimal trajectories for DMs, which could detrimentally impact both training duration and overall model performance. Fig. 9 shows that within $100$k iterations and for different numbers of function evaluations (NFE) after convergence, we consistently achieve a lower FID compared to the DM. In particular, the CFM model shows a faster reduction of the FID and provides better results.
Tab. 4 shows that the combination of DM and CFM outperforms the cascaded DMs across the board.
Taken together, these results underscore the training efficiency of our CFM model over DMs and its superior performance on the up-sampling task after fewer training steps.
Figure 9: Comparison of a diffusion-based [56] and our Coupling Flow Matching (CFM) module over the training for 4×4$\times$4 × up-sampling of the latent codes from $32^{2}\rightarrow 128^{2}$. The decoded output resolution is $1024^{2}$. We report FID and p-FID for a) different numbers of function evaluations (NFE) and b) throughout training. Architecture and hyperparameters are kept fixed. FID evaluated on $5$k samples from the LHQ validation set. We use $50$ steps for both DDIM [62] sampling and the Euler ODE solver.
Zero-shot LAION-5k 1024$\times$1024
$\bf{256^{2}\rightarrow 1024^{2}}$ $\bf{512^{2}\rightarrow 1024^{2}}$
Model Steps CLIP $\uparrow$ FID $\downarrow$ p-FID $\downarrow$ time (s/im) $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ p-FID $\downarrow$ time (s/im) $\downarrow$
Diffusion 4 22.92 35.22 50.53 0.19 25.68 26.53 25.20 2.72
CFM (ours) 4 23.77 27.54 24.02 0.19 26.16 21.61 15.83 2.72
Diffusion 40 23.55 26.67 24.16 0.62 26.05 22.29 16.36 3.16
CFM (ours) 40 23.75 25.47 23.31 0.62 26.14 21.67 15.96 3.16
Table 4: Quantitative comparison for $1024^{2}$ px image synthesis using SD1.5 [53] for sampling and either our Coupling Flow Matching (CFM) method or a diffusion-based latent space up-sampling model (DM) [56]. FID and Patch-FID are computed for 5k samples. We use the Euler ODE solver for CFM and DDIM sampling for the DM.
Figure 10: FID (left) and Patch-FID (right) for our model when applying different degrees of noise augmentation. Evaluated on $5$k samples.
Naïve Flow Matching. Finally, we compare with Naïve Flow Matching (FM). Similar to DM, FM is conditioned on the low-resolution latent code and starts with Gaussian noise, but uses an optimal transport-based objective to regress the vector fields. In contrast, our CFM method starts directly from the low-resolution latent code and regresses the vector field to the high-resolution counterpart. Due to the presence of data-dependent coupling, we have guaranteed optimal transport during training. For all methods, the low-resolution latent code is available as conditioning information throughout the entire generation trajectory. We evaluated the aforementioned two variants quantitatively (Tab. 3) and qualitatively (Fig. 8), where we observed that the CFM model with data-dependent coupling readily outperforms the ones without. We provide more information about the noise augmentation in Fig. 10. Notably, in the specific upsampling scenario from $256^{2}$ to $1024^{2}$, we observe an optimal configuration with a noising timestep of 400. The introduction of Gaussian noise proves beneficial as it imparts a smoothing effect on the input probability path, resulting in improved performance. However, excessive Gaussian noise can lead to the loss of valuable information, subsequently deteriorating the data-dependent coupling and reverting the model’s behavior to FM’s Gaussian assumption of $p(x_{0})$. This finding underscores the delicate balance required in incorporating noise for optimal model performance.
4.4 CFM for Degraded Image Super-Resolution
Our model is originally intended to render image synthesis with existing diffusion models more effectively by enabling them to operate on a lower resolution while increasing pixel-level resolution. However, our method can also be generalized to work on super-resolution tasks which usually include image degradations [67] for low-resolution images. By fine-tuning our method, we can achieve state-of-the-art results on two common benchmark datasets on a 4×4$\times$4 × upsampling task from $128^{2}$ to $512^{2}$ pixels. We provide quantitative (Tab. 5) and qualitative (Fig. 13) results in the appendix.
4.5 CFM Model Ablations
Upsampling Methods Since the dimensionality of the samples from both terminal distributions must be consistent for CFM, we need to upsample the low-resolution latent code $x_{0}$ to match the resolution at $x_{1}$. In this context, we perform an ablation study comparing three different upsampling methods: nearest neighbor upsampling, bilinear upsampling, and pixel space upsampling (PSU). The first two methods operate in latent space, while PSU requires the use of the KL autoencoder to upsample in pixel space. Denoting the latent encoder as $\mathcal{E}$, the decoder as $\mathcal{D}$, and the bilinear upsampling operation as UP, the upsampling operation PSU can be represented as $\mathcal{E}(\textit{\text{UP}}(\mathcal{D}(\cdot)))$. We empirically find that upsampling in latent space works well, but introduces artifacts that make distribution matching with CFM more difficult. In contrast, PSU yields faster model convergence at minimal additional cost (see Fig. 11) and also makes our approach invariant to the autoencoder used. Therefore, we use PSU unless otherwise stated.
Noise augmentation We systematically investigate the impact of varying levels of noise augmentation. Fig. 10 shows the FID and Patch-FID for different noise augmentation steps, with higher values corresponding to more noise. Our findings suggest that noise augmentation is crucial for model performance, albeit being quite robust to the amount of noise. Empirically, we discovered that $t=400$ yields the best results overall.
Intermediate Results along the ODE Trajectory In Fig. 14 we show intermediate results along the ODE trajectory. It can be seen that the CFM model gradually transforms the noise-augmented image representation to its high-resolution image counterpart.